# A finite element-based physics-informed operator learning framework for spatiotemporal partial differential equations on arbitrary domains

### v2

@@ Abstract
In this work, we propose a novel finite element-based physics-informed operator learning framework that allows for the prediction of dynamic physical phenomena governed by a spatiotemporal partial differential equation (PDE). A loss function inspired by the finite element method (FEM) with the implicit Euler time integration scheme is employed in the proposed framework. A transient thermal conduction problem is considered to benchmark the performance. The proposed operator learning framework takes a temperature field at the current time step as input and predicts a temperature field at the next time step. The Galerkin discretized weak formulation of the heat equation is employed to incorporate physics into the loss function, which is coined finite operator learning (FOL). Upon training on given boundary conditions and domain, the networks successfully predict the temperature evolution over time for any initial temperature field even after multiple time steps at high accuracy compared to the FEM solution. The framework is also confirmed to be applicable to cases with heterogeneous thermal conductivity, various time step sizes, and, more notably, arbitrary geometries. The advantages of the proposed FOL framework can be summarized as follows: First, the training of the networks is performed entirely in an unsupervised manner, which, in contrast to data-driven deep learning models, does not require the preparation of a large data set from costly simulations or experiments. Instead, a data set of random temperature patterns generated by the Gaussian random process and the Fourier series, combined with constant temperature fields, is used as training data to cover a wide range of possible temperature cases. In addition, shape functions and backward difference approximation are exploited for spatial and temporal discretization, resulting in a purely algebraic equation. This enhances training efficiency, as one avoids time-consuming automatic differentiation when optimizing weights and biases, although we should anticipate and accept possible errors arising from the discretization of the physical domain. Finally, thanks to the interpolation power of the finite element approach, any arbitrary geometry can be handled with FOL, which is crucial to addressing various scenarios in engineering applications. The potential of the framework for other spatiotemporal PDEs is also discussed.
@@ 1 Introduction
Over the past decade, machine learning (ML) methods have played a prominent role in scientific and engineering applications. They can learn how to perform a given task that previously could only be done by humans. The famous applications include self-driving cars [1], natural language processing [2], and image recognition [3]. Many of these applications utilize deep learning (DL), a subset of ML that has attracted attention for its ability to perform the mapping between input and output features. Accurate prediction by DL can be achieved by sufficiently training artificial neural networks (NNs). When it comes to the field of computational mechanics, there are various types of boundary value problems (BVP) described by partial differential equations (PDEs), which are commonly solved using numerical methods such as the finite element method (FEM), finite difference method, mesh-free method, etc. The problems with the use of numerical schemes can be the computational cost, the complexity of the mesh generation, and, more importantly, the fact that each simulation has to be performed almost from scratch for every scenario [4]. DL has been utilized as a promising alternative to avoid these problems. Two mainstreams exist in the domain of DL applications to computational mechanical problems. One is supervised learning based on available labeled data. For example, Liang et al. demonstrate the potential of a DL model to be a surrogate of FE analysis in estimating the stress distribution of cardiovascular vessels, which allows for fast and accurate predictions of stress distribution in biomedical applications [5]. Hsu et al. present a DL-based predictive model for crack propagation of crystalline materials using image data processed from the visualized results of molecular dynamics (MD) simulations [6]. Furthermore, Fernandez et al. propose a DL model on the constitutive behavior of grain boundaries, which takes the traction-separation effects into account, based on the data obtained from MD simulations [7]. Many other DL models on computational mechanical problems have also been developed within the scope of supervised learning; see [8, 9, 10, 11, 12, 13] as examples.
@@
The other common approach is unsupervised learning based on governing equations of BVPs, even without labeled data for training. Originally proposed by Raissi et al., NNs trained based on physics-based loss functions from BVPs are called physics-informed neural networks (PINNs) [14]. The key idea is to incorporate governing PDEs directly into the loss functions of NNs with the power of automatic differentiation. Upon successful training, PINNs can accurately predict physical behaviors within the domain of a problem. The training can be seen as the minimization problem in which the residual of PDEs is used as a target function. For the last five years, many researchers have tested the capability of PINNs to predict the behavior of a physical system. Jin et al. developed a PINN framework for incompressible Navier-Stokes equations and verified its capability of obtaining approximate solutions to ill-posed problems with noisy boundary conditions and inverse problems in the context of flow simulation [15]. Mao et al. modeled high-speed flow based on the Euler equation using PINNs [16]. Mahmoudabadbozchelou et al. presented non-Newtonian PINNs for solving coupled PDEs for fluid while considering the constitutive relationships [17]. Besides, many other studies on fluid-oriented applications of PINNs, such as [18, 19, 20, 21], have already been investigated in recent years. For heat conduction problems, Zobeiry et al. applied the PINN architecture to the heat transfer equation with convective boundary conditions [22]. Cai et al. modeled heat convections with unknown boundary conditions and the two-phase Stefan problem [23]. Zhao et al. developed a combined framework of PINNs and convolutional neural networks (CNNs) for predictions of temperatures from the information of heat source [24]. Furthermore, Guo et al. worked on the prediction of three-dimensional transient heat conduction targeted for functionally graded materials using the deep collocation method for space and the Runge-Kutta scheme for time integration, showing the applicability of PINN approaches to spatiotemporal three-dimensional complex geometry cases [25]. Readers can also refer to [26, 27, 28, 29, 30] for other PINNs examples on heat transfer problems. When it comes to solid mechanics problems, Samaniego et al. developed a variational energy-based physics-informed loss function for the classical linear elasticity problem and the phase-field model for fracture [31]. Abueidda et al. used the collocation method to solve solid mechanics problems with various types of material models, including hyperelasticity with large deformation [32]. Haghighat et al. demonstrated the applicability of PINNs to the von Mises plasticity model in their PINN framework [33]. Rezaei et al. proposed a PINN solver for solid problems with heterogeneous elasticities [34]. Harandi et al. solved the thermomechanical coupled system of equations in the heterogeneous domains [35]. Bai et al. developed a modified loss function using the least squares weighted residual method for two- and three-dimensional solid mechanics, which can predict well the displacement and stress fields [36]. Other investigations into PINNs for solid mechanics can also be found in [37, 38].
@@
Another open problem in physics-informed deep learning is the failure to predict time-dependent evolutionary processes. Wang et al. argued that the causality of physics must be respected in training PINNs when one considers time-continuous problems [56]. This is the case when we directly treat the temporal dimension as an additional dimension to the spatial domain. Mattey et al. developed a PINN model that enforces backward compatibility over the temporal domain in the loss function to overcome this limitation in the Allen-Cahn and Cahn-Hilliard equations [57]. Xu et al. utilized transfer learning for DeepONet to train the networks with better stability than the original DeepONet for dynamic systems [41]. Li et al. presented a phase-field DeepONet, which aims to predict the dynamic behavior of phase-fields in the Allen-Cahn and Cahn-Hilliard equations using the concept of gradient flows [58]. In the latter framework, the trained networks work as an explicit time-stepper that can predict the evolution of the phase field at the next step based on the current phase field. Furthermore, an emerging approach for spatiotemporal predictions is the utilization of numerical discretizations or convolutions to discretize derivatives to learn a discrete mapping on a discretized domain. This direction can also enhance training efficiency by avoiding time-consuming automatic differentiation, especially when higher-order derivatives need to be computed. For static problems, Fuhg et al. proposed a deep convolutional Ritz method as a surrogate of numerical solvers, in which the convolution is exploited to take central differences, and the network takes the energy form as a physics-informed loss[59]. Gao et al. utilized CNN architecture to deal with the discretized domain and extended it to irregular domains through coordinate transformation [60]. Rezaei et al. devised a framework that they named finite operator learning (FOL) based on FEM for parametrically solving PDEs with a demonstration for a steady heat equation with heterogeneity [61]. Some other works have applied FEM to integrate the weak-form loss into NNs, such as for advection-diffusion [62], quantification of wind effects on vibrations [63], etc. When it comes to spatiotemporal problems, Geneva et al. proposed a CNN-based framework with autoregressive encoder-decoder architecture, whose performance is showcased for some types of dynamic PDEs [64]. Ren et al. presented a discrete learning architecture that combines CNN with long short-term memory for spatiotemporal PDEs [65]. Liu et al. embedded known PDE information into CNN architecture itself to preserve the behavior of the PDE of interest for spatiotemporal dynamic phenomena [66]. Furthermore, Xiang et al. employed graph neural networks in combination with radial basis function finite difference to predict spatiotemporal dynamics for irregular domains [67]. The abovementioned works have shown the capability of discrete mapping learned by NNs for spatiotemporal dynamics. However, the researchers in this domain are still looking for approaches that can easily address irregular domains, as it is difficult for CNN-based methods in particular. In this sense, the direction of the incorporation of FEM into discrete operator learning for parametrically solving spatiotemporal PDEs is beneficial to address more realistic problem setups.
@@
In this study, we aim to develop a novel physics-informed discrete-type operator learning framework, which we refer to as FOL, that can parametrically predict the dynamic behavior of physical quantities over time. The schematic illustration of the developed framework is provided in Fig. 1. The key idea is to provide physical fields of a system at the current time step as input and return those at the next step as output, realizing a surrogate model of time-marching numerical schemes. The time-dependent heat equation, also known as a transient heat conduction problem, is chosen as the target BVP to validate the framework proposed in this work. Not only does this study consider homogeneous thermal conductivity, but it also takes into account heterogeneous conductivity. The training follows a physics-informed loss function constructed based on the finite element discretization of the heat equation [61], thereby making it unsupervised learning without labeled data. The extension of the framework to irregular domains is also tested at the end.
@@
The difference from the previous frameworks, such as the one by [58] or by [66] for example, is that this framework directly uses the discretized weak form loss that is identical to the formulation when solving with FEM. This is also demonstrated in a representative model later in this paper. Furthermore, this study considers the heterogeneity of physical properties, which is not addressed in the aforementioned studies.
@@ 3.2 Proposed finite element-based physics-informed operator learning framework
The core idea of the FOL framework is to predict physical fields at the next time step, utilizing their current time step state, which is equivalent to other time marching FE solvers. The network architecture is shown in Fig. 5, implemented using the TensorFlow-based deep learning library SciANN [33]. The domain is first discretized through finite elements; see Fig. 6. The nodes in the discretized domain are the representative points for evaluating temperature evolution by the networks. Analogous to the finite element method, the Gaussian integration is performed to integrate over the elements using the bi-linear shape function, shown in the right of Fig. 6. Regarding the network architecture, it is worth noting that separate feedforward NNs are used to predict each node’s temperature output. In [61], the authors showed that separate networks with a small number of neurons per layer in each network outperformed a single fully connected network with a large number of neurons per layer. All the separated NNs are trained at the same time through a physically informed loss function based on the input and output temperature fields. Substituting , which means backward Euler approximation in time, into Eq. 7 and taking the L2 norm of the residual yields the loss function in this framework, which reads,
@@ 4 Results
Using the framework introduced in the previous section, we have attempted to predict the solution of the heat equation for any given initial temperature field. As a demonstration, we show here only the results of the predictions for three of the initial temperature fields, one with , one from the Gaussian distribution, and the last one with , considered in this work. We can confirm that the chosen initial temperatures are not exactly represented in the initial training samples. Furthermore, as will be shown later, the temporal temperature evolution results in complex patterns that significantly differ from the training samples. It is important to note that since the training is entirely unsupervised, the network is not provided with solutions for these initial temperature samples. Hence, the subsequent results serve as rigorous tests to evaluate the network’s performance. In the main study, we trained the NNs for epochs to ensure that the loss reaches a plateau due to convergence, which can be confirmed in Fig. 8, although it continues to decrease slightly over the epoch.
@@
In post-processing, we calculated the heat flux based on the obtained temperature fields with respect to the discretized system to further investigate the physical behavior of the predicted heat conduction. When the homogeneous thermal conductivity is applied, Fig. 9 exhibits that the FOL predictions agree with the reference FE solution with a maximum error of about in the nodal absolute temperature error and in the nodal absolute heat flux magnitude error at [s] with as the initial temperature field, where in Fig. 9 and denote the difference in temperature and heat flux magnitude between the prediction and reference solution at each node, respectively. The heat flux transition over time shows that the transient behavior is accurately predicted by the proposed FOL framework upon training. The absolute error distribution shown at the bottom of Fig. 9 looks rather random. The error can be further reduced by, for example, enhancing the variety of training samples, which will lead to better coverage of possible temperature fields in the network input. Looking at the results in Fig. 10, we confirmed that even for a random temperature pattern generated from the Gaussian distribution, one can obtain reasonable predictions with a maximum of in nodal absolute temperature error and in nodal absolute heat flux magnitude error at [s]. In Fig. 11, the temperature error increased by a factor of to approximately at maximum in the nodal absolute temperature error, while the heat flux magnitude error increased by a factor of to about . In Fig. 15 in the nodal absolute heat flux magnitude error at [s] with as the initial temperature field. Nevertheless, the prediction correctly captured the main feature of the temperature evolution seen in the reference solution by FEM in the qualitative comparison of the results from FOL and FEM. When it comes to the case with the heterogeneous thermal conductivity, one can also see in Figs. 12, 13 and 14 the agreement of the FOL prediction with the corresponding FEM solutions. The main difference to the homogeneous case is that the error accumulation in the temperature field is dominant around the inserted low conductivity regions (i.e., phase boundaries) due to steep changes in the solution. To enhance the solution’s quality, one can train the neural network with additional input neurons and utilize advanced optimizers, such as L-BFGS, with hyperparameter tuning to prevent getting trapped in local minima. It is important to note that the prediction of dynamic temperature evolution in the other part of the domain is reasonable. One can also confirm in Figs. 12, 13, and 14 that the red arrows representing the heat flux bypass the low conductivity region area in both FOL and FEM. Furthermore, we also predicted for more time steps up to [s] to see if a long-term prediction is possible; see Fig. 15. The prediction was still in good agreement with the FE solution even at , although error concentration is seen in the low conductivity region. Overall, the results of the demonstration on the transient heat conduction problem show that the proposed FOL framework for spatiotemporal PDEs is capable of predicting the solution and even its spatial gradients under a given time step size and boundary conditions.

### v3

@@ Abstract
We propose a novel finite element-based physics-informed operator learning framework that allows for predicting spatiotemporal dynamics governed by partial differential equations (PDEs). The Galerkin discretized weak formulation is employed to incorporate physics into the loss function, termed finite operator learning (FOL), along with the implicit Euler time integration scheme for temporal discretization. A transient thermal conduction problem is considered to benchmark the performance, where FOL takes a temperature field at the current time step as input and predicts a temperature field at the next time step. Upon training, the network successfully predicts the temperature evolution over time for any initial temperature field at high accuracy compared to the solution by the finite element method (FEM) even with a heterogeneous thermal conductivity and arbitrary geometry. The advantages of FOL can be summarized as follows: First, the training is performed in an unsupervised manner, avoiding the need for large data prepared from costly simulations or experiments. Instead, random temperature patterns generated by the Gaussian random process and the Fourier series, combined with constant temperature fields, are used as training data to cover possible temperature cases. Additionally, shape functions and backward difference approximation are exploited for the domain discretization, resulting in a purely algebraic equation. This enhances training efficiency, as one avoids time-consuming automatic differentiation in optimizing weights and biases while accepting possible discretization errors. Finally, thanks to the interpolation power of FEM, any arbitrary geometry with heterogeneous microstructure can be handled with FOL, which is crucial to addressing various engineering application scenarios.
@@ 1 Introduction
Over the past decade, machine learning (ML) methods have played a prominent role in scientific and engineering applications. They can learn how to perform a given task that previously could only be done by humans. The famous applications include self-driving cars [1], natural language processing [2], and image recognition [3]. Many of these applications utilize deep learning (DL), a subset of ML that has attracted attention for its ability to perform the mapping between input and output features. Accurate prediction by DL can be achieved by sufficiently training artificial neural networks (NNs).
@@
When it comes to the field of computational mechanics, there are various types of boundary value problems (BVP) described by partial differential equations (PDEs), which are commonly solved using numerical methods such as the finite element method (FEM), finite difference method, mesh-free method, etc. The problems with the use of numerical schemes can be the computational cost, the complexity of the mesh generation, and, more importantly, the fact that each simulation has to be performed almost from scratch for every scenario [4]. DL has been utilized as a promising alternative to avoid these problems. Two mainstreams exist in the domain of DL applications to computational mechanical problems. One is supervised learning based on available labeled data. For example, Liang et al. demonstrated the potential of a DL model to be a surrogate of FE analysis in estimating the stress distribution of cardiovascular vessels, which allows for fast and accurate predictions of stress distribution in biomedical applications [5]. DL for biomedical applications has also been seen in the prediction of adolescent idiopathic scoliosis [6] and pediatric spinal deformities [7], in which X-ray data is used as clinical input and the results of calculations by FEM are employed as mechanistic input. One can also see the employment of graph neural networks for the prediction of material concentration in neurite networks [8] and the combination of the isogeometric analysis with convolutional neural networks (CNNs) for the prediction of neuron growth [9]. In addition, Li et al. proposed an encoder-decoder-based CNN for reaction-diffusion systems as a fast and accurate surrogate tool to FEM [10]. In material modeling, Hsu et al. presented a DL-based predictive model for crack propagation of crystalline materials using image data processed from the visualized results of molecular dynamics (MD) simulations [11]. Furthermore, Fernandez et al. proposed a DL model on the constitutive behavior of grain boundaries, which takes the traction-separation effects into account, based on the data obtained from MD simulations [12]. Studies have also been conducted that delve into learning differential operators for PDEs from data [13, 14]. Many other DL models on computational mechanical problems have also been developed within the scope of supervised learning; see [15, 16, 17, 18, 19, 20] as examples.
@@
The other common approach is unsupervised learning based on governing equations of BVPs, even without labeled data for training. Originally proposed by Raissi et al., NNs trained based on physics-based loss functions from BVPs are called physics-informed neural networks (PINNs) [21]. The key idea is to incorporate governing PDEs directly into the loss functions of NNs with the power of automatic differentiation. Upon successful training, PINNs can accurately predict physical behaviors within the domain of a problem. The training can be seen as the minimization problem in which the residual of PDEs is used as a target function. For the last five years, many researchers have tested the capability of PINNs to predict the behavior of a physical system. Jin et al. developed a PINN framework for incompressible Navier-Stokes equations and verified its capability of obtaining approximate solutions to ill-posed problems with noisy boundary conditions and inverse problems in the context of flow simulation [22]. Mao et al. modeled high-speed flow based on the Euler equation using PINNs [23]. Mahmoudabadbozchelou et al. presented non-Newtonian PINNs for solving coupled PDEs for fluid while considering the constitutive relationships [24]. Besides, many other studies on fluid-oriented applications of PINNs, such as [25, 26, 27, 28], have already been investigated in recent years. For heat conduction problems, Zobeiry et al. applied the PINN architecture to the heat transfer equation with convective boundary conditions [29]. Cai et al. modeled heat convections with unknown boundary conditions and the two-phase Stefan problem [30]. Zhao et al. developed a combined framework of PINNs and CNNs for predictions of temperatures from the information of heat source [31]. Furthermore, Guo et al. worked on the prediction of three-dimensional transient heat conduction targeted for functionally graded materials using the deep collocation method for space and the Runge-Kutta scheme for time integration, showing the applicability of PINN approaches to spatiotemporal three-dimensional complex geometry cases [32]. Readers can also refer to [33, 34, 35, 36, 37] for other PINNs examples on heat transfer problems. When it comes to solid mechanics problems, Samaniego et al. developed a variational energy-based physics-informed loss function for the classical linear elasticity problem and the phase-field model for fracture [38]. Abueidda et al. used the collocation method to solve solid mechanics problems with various types of material models, including hyperelasticity with large deformation [39]. Haghighat et al. demonstrated the applicability of PINNs to the von Mises plasticity model in their PINN framework [40]. Rezaei et al. proposed a PINN solver for solid problems with heterogeneous elasticities [41]. Harandi et al. solved the thermomechanical coupled system of equations in the heterogeneous domains [42]. Bai et al. developed a modified loss function using the least squares weighted residual method for two- and three-dimensional solid mechanics, which can predict well the displacement and stress fields [43]. Other investigations into PINNs for solid mechanics can also be found in [44, 45]. In addition, the idea of PINNs has also been combined with the isogeometric analysis for predicting material transports in neurons [46].
@@
Another open problem in physics-informed deep learning is the failure to predict time-dependent evolutionary processes. Wang et al. argued that the causality of physics must be respected in training PINNs when one considers time-continuous problems [64]. This is the case when we directly treat the temporal dimension as an additional dimension to the spatial domain. Mattey et al. developed a PINN model that enforces backward compatibility over the temporal domain in the loss function to overcome this limitation in the Allen-Cahn and Cahn-Hilliard equations [65]. Xu et al. utilized transfer learning for DeepONet to train the networks with better stability than the original DeepONet for dynamic systems [49]. Li et al. presented a phase-field DeepONet, which aims to predict the dynamic behavior of phase-fields in the Allen-Cahn and Cahn-Hilliard equations using the concept of gradient flows [66]. In the latter framework, the trained networks work as an explicit time-stepper that can predict the evolution of the phase field at the next step based on the current phase field. Furthermore, an emerging approach for spatiotemporal predictions is the utilization of numerical discretizations or convolutions to discretize derivatives to learn a discrete mapping on a discretized domain. This direction can also enhance training efficiency by avoiding time-consuming automatic differentiation, especially when higher-order derivatives need to be computed. For static problems, Fuhg et al. proposed a deep convolutional Ritz method as a surrogate of numerical solvers, in which the convolution is exploited to take central differences, and the network takes the energy form as a physics-informed loss [67]. Gao et al. utilized CNN architecture to deal with the discretized domain and extended it to irregular domains through coordinate transformation [68]. Rezaei et al. devised a framework that they named finite operator learning (FOL) based on FEM for parametrically solving PDEs with a demonstration for a steady heat equation with heterogeneity [69]. Some other works have applied FEM to integrate the weak-form loss into NNs, such as for advection-diffusion [70], quantification of wind effects on vibrations [71], etc. Furthermore, Khara et al. employed the energy-form loss in the FEM-inspired loss function and demonstrated its performance in Poisson’s equation including a three-dimensional case [72]. When it comes to spatiotemporal problems, Geneva et al. proposed a CNN-based framework with autoregressive encoder-decoder architecture, whose performance is showcased for some types of dynamic PDEs [73]. Ren et al. presented a discrete learning architecture that combines CNN with long short-term memory for spatiotemporal PDEs [74]. Liu et al. embedded known PDE information into CNN architecture itself to preserve the behavior of the PDE of interest for spatiotemporal dynamic phenomena [75]. Furthermore, Xiang et al. employed graph neural networks in combination with radial basis function finite difference to predict spatiotemporal dynamics for irregular domains [76]. The abovementioned works have shown the capability of discrete mapping learned by NNs for spatiotemporal dynamics. However, the researchers in this domain are still looking for approaches that can easily address irregular domains, as it is difficult for CNN-based methods in particular. In this sense, the direction of the incorporation of FEM into discrete operator learning for parametrically solving spatiotemporal PDEs is beneficial to address more realistic problem setups.
@@
In this study, we aim to develop a novel physics-informed discrete-type operator learning framework, which we refer to as FOL, that can parametrically predict the dynamic behavior of physical quantities over time. The schematic illustration of the developed framework is provided in Fig. 1. The key idea is to provide physical fields of a system at the current time step as input and return those at the next time step as output, realizing a surrogate model of time-marching numerical schemes. The time-dependent heat equation, also known as a transient heat conduction problem, is chosen as the target BVP to validate the framework proposed in this work. Not only does this study consider homogeneous thermal conductivity, but it also takes into account heterogeneous conductivity. The training follows a physics-informed loss function constructed based on the finite element discretization of the heat equation [69], thereby making it unsupervised learning without labeled data. The extension of the framework to irregular domains is also tested at the end.
@@
The difference from the previous frameworks, such as the one by [66] or by [75] for example, is that this framework directly uses the discretized weak form loss that is identical to the formulation when solving with FEM. This is also demonstrated in a representative model later in this paper. Furthermore, this study considers the heterogeneity of physical properties, which is not addressed in the aforementioned studies. For clarity, the comparison of the architecture with the vanilla PINNs and physics-informed DeepONet (PI-DeepONet) is described in Fig. 2. The pivotal difference is that in FOL we embed the coordinate information into the loss function, allowing us to integrate the branch and trunk nets in DeepONet into a single network. In addition, we do not take the temporal coordinate as input unlike PINNs or PI-DeepOnet; FOL takes into account the temporal evolution by discretizing a given PDE in time between the current and next time steps.
@@ 3.2 Proposed finite element-based physics-informed operator learning framework
The core idea of the FOL framework is to predict physical fields at the next time step, utilizing their current time step state, which is equivalent to other time marching FE solvers. The network architecture is shown in Fig. 6, implemented using the TensorFlow-based deep learning library SciANN [40]. The domain is first discretized through finite elements; see Fig. 7. The nodes in the discretized domain are the representative points for evaluating temperature evolution by the networks. Analogous to the finite element method, the Gaussian integration is performed to integrate over the elements using the bi-linear shape function, shown in the right of Fig. 7. Regarding the network architecture, it is worth noting that separate feedforward NNs are used to predict each node’s temperature output. In [69], the authors showed that separate networks with a small number of neurons per layer in each network outperformed a single fully connected network with a large number of neurons per layer. Nevertheless, it is also shown in the same work that using a simple fully connected network with a properly reduced architecture performs very well in finding the correct solutions. Therefore, the user needs to study this matter according to the problem at hand and the nature of the equations and outputs. The comparison of the performance between the separated network architecture and the fully connected architecture is described in Section 4.5. All the separated NNs are trained at the same time through a physically informed loss function based on the input and output temperature fields. Substituting , which means backward Euler approximation in time, into Eq. 7 and taking the L2 norm of the residual yields the loss function in this framework, which reads,
@@ 4 Results
Using the framework introduced in the previous section, we have attempted to predict the solution of the heat equation for any given initial temperature field. As a demonstration, we show here only the results of the predictions for three of the initial temperature fields, one with , one from the Gaussian distribution, and the last one with , considered in this work. We can confirm that the chosen initial temperatures are not exactly represented in the initial training samples. Furthermore, as will be shown later, the temporal temperature evolution results in complex patterns that significantly differ from the training samples. It is important to note that since the training is entirely unsupervised, the network is not provided with solutions for these initial temperature samples. Hence, the subsequent results serve as rigorous tests to evaluate the network’s performance. In the main study, we trained the NNs for epochs to ensure that the loss reaches a plateau due to convergence after hundreds of epochs. This is confirmed in Fig. 9, although it continues to decrease slightly over the epoch. We do not employ fixed criteria on the residual value for stopping the training, as this is currently not the main scope of this work.
@@
In post-processing, we calculated the heat flux based on the obtained temperature fields with respect to the discretized system to further investigate the physical behavior of the predicted heat conduction. When the homogeneous thermal conductivity is applied, Fig. 10 exhibits that the FOL predictions agree with the reference FE solution with a maximum error of about in the nodal absolute temperature error and in the nodal absolute heat flux magnitude error at [s] with as the initial temperature field, where in Fig. 10 and denote the difference in temperature and heat flux magnitude between the prediction and reference solution at each node, respectively. The heat flux transition over time shows that the transient behavior is accurately predicted by the proposed FOL framework upon training. The absolute error distribution shown at the bottom of Fig. 10 looks rather random. The error can be further reduced by, for example, enhancing the variety of training samples, which will lead to better coverage of possible temperature fields in the network input. Looking at the results in Fig. 11, we confirmed that even for a random temperature pattern generated from the Gaussian distribution, one can obtain reasonable predictions with a maximum of in nodal absolute temperature error and in nodal absolute heat flux magnitude error at [s]. Here, we define ”reasonable prediction” as the relative L2 error being small against the finite element solution, which means the predictions by FOL capture the important features of the temperature evolutions. More concretely, for , if the relative L2 error in temperature prediction is less than 0.1, the FOL prediction is at reasonable accuracy for example. In Fig. 12, the temperature error increased by a factor of to approximately at maximum in the nodal absolute temperature error, while the heat flux magnitude error increased by a factor of to about . In Fig. 15 in the nodal absolute heat flux magnitude error at [s] with as the initial temperature field. Nevertheless, the prediction correctly captured the main feature of the temperature evolution seen in the reference solution by FEM in the qualitative comparison of the results from FOL and FEM. When it comes to the case with the heterogeneous thermal conductivity, one can also see in Figs. 13, 14 and 15 the agreement of the FOL prediction with the corresponding FEM solutions. The main difference to the homogeneous case is that the error accumulation in the temperature field is dominant around the inserted low conductivity regions (i.e., phase boundaries) due to steep changes in the solution. To enhance the solution’s quality, one can train the neural network with additional input neurons and utilize advanced optimizers, such as L-BFGS, with hyperparameter tuning to prevent getting trapped in local minima. It is important to note that the prediction of dynamic temperature evolution in the other part of the domain is reasonable. One can also confirm in Figs. 13, 14, and 15 that the red arrows representing the heat flux bypass the low conductivity region area in both FOL and FEM. Furthermore, we also predicted for more time steps up to [s] to see if a long-term prediction is possible; see Fig. 16. The prediction was still in good agreement with the FE solution even at , although error concentration is seen in the low conductivity region. Overall, the results of the demonstration on the transient heat conduction problem show that the proposed FOL framework for spatiotemporal PDEs is capable of predicting the solution and even its spatial gradients under a given time step size and boundary conditions.